OpenStudy (drewkatski):

My teacher and his worksheets from online say that the formula for Simple Interest is [I = PRT/100] but wherever I look online, I see its just [PxRxT]... Which one is it? -_-

Directrix (directrix):

Look at the attached example. R for rate is expressed as a percentage. --- http://www.mathsisfun.com/money/interest.html

Directrix (directrix):

It appears that your teacher's simple interest formula is the same. [I = PRT/100] with the rate given not as a percentage. Formulas don't mean much unless the variables in the formulas are defined. @Drewkatski Do you have an interest problem you would like to work? If so, post it. Thanks.

OpenStudy (drewkatski):

Ok, so its the same :) I understand how to do Simple Interest, I just wondered if my teacher was giving me the right thing or not

OpenStudy (drewkatski):

But why exactly am I dividing the PRT answer by 100?

Directrix (directrix):

Your teacher gives the problem: Jan borrowed $43,000 for 4 years at a 5% interest rate. How much interest is that? Teacher's Formula I = (P*R*T) / 100 where P = $3,000, R = 5, and T = 4 years Not that R is 5 in the formula. Yet, the interest rate is 5%. Percent means hundreths. So 5% = 5/100. Rather than have you make that calculation, the teacher put the 100 in the denominator so that the 5 in the formula is really 5/100 = 5%.

Directrix (directrix):

In the Internet problem I posted, the person who did that problem used the formula I = PRT where P is principal, R = interest percentage expressed as a decimal, and T is the time in years of the loan. In short, the two formulas are the same. My suggestion is that you learn one and stick with it.

OpenStudy (drewkatski):

Ok, I will go with R = PRT, it'll save time :D Thanks

Directrix (directrix):

Okay, but remember the variables: I for simple interest amount, P for principal, R for interest rate expressed as a decimal, and T for time in years. @Drewkatski

OpenStudy (drewkatski):

yep